OpenStudy (anonymous):

JCL Computers has five vacancies in its executive trainee program. In how many ways can the company select five trainees from a group of fifteen female and fifteen male applicants if the vacancies (b) Must be filled by two men and three women?

OpenStudy (anonymous):

you are being asked how many ways to choose 2 out of 15 and 3 out of 15 do you know how to do that?

OpenStudy (anonymous):

I thought it was 15*14 + 15*14*13 but its not

OpenStudy (anonymous):

i.e. you are being asked for \[\binom{15}{2}\times \binom{15}{3}\] some times written as \[_{15}C_2\times _{15}C_3\]

OpenStudy (anonymous):

ok , it should be times, not plus, by the "counting principle" so one mistake the the plus sign

OpenStudy (anonymous):

but another mistake is \[\binom{15}{2}=\frac{15\times 14}{2}\] and \[\binom{15}{3}=\frac{15\times 14\times 13}{3\times 2}\]

OpenStudy (anonymous):

and again, it is times, not plus

OpenStudy (anonymous):

if it was me, i would cheat rather than bother with a calculator or arithmetic http://www.wolframalpha.com/input/?i=%2815+choose+2%29%2815+choose+3%29

OpenStudy (anonymous):

thank you!!!